Start learning →

For an event $E$ and its complement $\overline{E}$ (not E),

A$P(E) + P(\overline{E}) = 0$
B$P(E) = P(\overline{E})$
C$P(E) + P(\overline{E}) = 1$
D$P(E) - P(\overline{E}) = 1$
Answer & Solution
Correct answer: C. $P(E) + P(\overline{E}) = 1$
1. Either the event happens or it does not. 2. So $P(E) + P(\overline{E}) = 1$, i.e. $P(\overline{E}) = 1 - P(E)$. 3. This is the complement rule. _Source: Karnataka SSLC (KSEEB) Class 10 Mathematics, Ch14 'Probability', §14.1_
Practise this question with Quill — Karnataka SSLC (Class 10) →

Keep practising Probability

Browse all Probability questions for Karnataka SSLC (Class 10), or explore more Mathematics topics.

Start a Probability lesson →
Related questions